OpenStudy (anonymous):

Solve Equations: 2(x+3)-1=7x

OpenStudy (anonymous):

distribute the 2

OpenStudy (anonymous):

so 2x + 6 -1 = 7x

OpenStudy (anonymous):

Step 1) Distribute Step 2) Isolate variable terms Step 3) Combine terms Step 4) Divide by the coefficient on the variable

OpenStudy (anonymous):

6-1 = 5

OpenStudy (anonymous):

so then 2x + 5= 7x -2x -2x _________________________ 5= 5x x=1

OpenStudy (anonymous):

does that help?